Part A Find the capacitance of a parallel-plate capacitor consisting of circular plates 34 cm in radius separated by 2.4 mm . Express your answer using two significant figures.

C = in pF
Answer #1

According to the question we need to find the value of the capacitance.

r = 34.0 cm = 0.34 m d = 2.40 mm Area= 3.14 * ( 0.34 * 0.34 m ) = 0.362904 m^2

C = A * ( 8.85 * 10^ -12 ) / ( 2.40 * 10 ^ -3 m ) = 1338.5035 pF ======================ANSWER)